OpenStudy (anonymous):

linear equations Kelly ordered a sub that cost 4.95 and additional 0.35 per topping. If kellys bill for the sub is 6.70 how many additional toppings did she add

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

\[\bf \large 4.95+.35x=6.7\] Solve for x

OpenStudy (anonymous):

one sec then

OpenStudy (anonymous):

-4.96-6.7=11.66 right brb

Vocaloid (vocaloid):

uh, not quite, you want to subtract 4.95 from each side... left side: 4.95+.35x right side: 6.7 4.95+.35x - 4.95 = 6.7 - 4.95 keep going from here

OpenStudy (anonymous):

-1.74

Vocaloid (vocaloid):

4.95+.35x - 4.95 = ?

OpenStudy (anonymous):

9.9

Vocaloid (vocaloid):

not quite, remember that you can't combine variables and numbers... 4.95 - 4.95 = ?

OpenStudy (anonymous):

0

Vocaloid (vocaloid):

right, so 4.95+.35x - 4.95 = ?

OpenStudy (anonymous):

35x=0

Vocaloid (vocaloid):

4.95 - 4.95 = 0 4.95 + 0.35x - 4.95 = 0.35x, understood? we subtract 4.95 - 4.95 and leave 0.35x alone since it's got the letter x attached to it

Vocaloid (vocaloid):

now let's look at the right side: 6.7 - 4.95 = ?

OpenStudy (anonymous):

1.74

Vocaloid (vocaloid):

check your math again

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

1.45

Vocaloid (vocaloid):

check it again...

Vocaloid (vocaloid):

use a calculator if you are unsure

OpenStudy (anonymous):

i am

Vocaloid (vocaloid):

it's not correct, keep trying

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i'm getting -1.74

Vocaloid (vocaloid):

type 6.7 - 4.95 into a calculator

OpenStudy (anonymous):

1.75

Vocaloid (vocaloid):

right. so 0.35x = 1.75 solve for x

OpenStudy (anonymous):

you know i had the same answer up there

Vocaloid (vocaloid):

no you didn't...

Vocaloid (vocaloid):

your last two answers were -1.74 and 1.45....

OpenStudy (anonymous):

oh it was 1.74 right

Vocaloid (vocaloid):

so 0.35x = 1.75 solve for x

OpenStudy (anonymous):

1.4

Vocaloid (vocaloid):

so 0.35x = 1.75 divide both sides by 0.35...

OpenStudy (anonymous):

ohh right sorry

OpenStudy (anonymous):

5

Vocaloid (vocaloid):

yup, that's our final answer

OpenStudy (anonymous):

thanks. :)